# Diffuse Light in Milky-Way like Haloes

### 26 Jan 2024 v1

@@ Abstract
We investigate the diffuse light (DL) content of dark matter haloes in the mass range $11.5\leq\log M_{halo}\leq 13$, a range that includes also the dark matter halo of the Milky-Way, taking advantage of a state-of-the-art semi-analytic model run on the merger trees extracted from a set of high-resolution cosmological simulations. The fraction of DL in such relatively small haloes is found to progressively decrease from the high to the low mass end, in good agreement with analytic (Purcell et al. 2007) and numerical results from simulations (Proctor et al. 2023; Ahvazi et al. 2023), in good agreement also with the fraction of the DL observed in the Milky-Way (Deason et al. 2019) and M31 (Harmsen et al. 2017). Haloes with different masses have a different efficiency in producing DL: $\log M_{halo}\simeq 13$ is found to be the characteristic halo mass where the production of DL is the most efficient, while the overall efficiency decreases at both larger (Contini et al. 2024) and smaller scales (this work). The DL content in this range of halo mass is the result of stellar stripping due to tidal interaction between satellites and its host (95%) and mergers between satellites and the central galaxy (5%), with pre-processed material, sub-channel of mergers and stripping and so already included in the 100%, that contributes no more than 8% on average. The halo concentration is the main driver of the DL formation: more concentrated haloes have higher DL fractions that come from stripping of more massive satellites in the high halo mass end, while dwarfs contribute mostly in the low halo mass end.
@@ 3 Results
Most of the DL forms through stellar stripping, and we already know (Contini et al. 2014) that intermediate/massive galaxies are responsible for the bulk of it. The last question we want to answer, in order to complete the analysis is: are intermediate/massive galaxies still the main contributors? The answer to this question is in Figure 5, where we plot the percentage of DL that comes from stellar stripping of galaxies having a given stellar mass (top panel), for haloes in different mass ranges as indicated in the legend of the Figure. We can clearly see that the peak of the distributions progressively decreases, from $\log M_{*}\sim 10.3$ in the highest halo mass range, to $\log M_{*}\sim 9.0$ in the lowest. The contribution of galaxies in the stellar mass range $11.09.5<\log M_{*}<$ 11.0 increases from 16% to 62% in the same ranges, implying that while intermediate/massive galaxies still contribute remarkably in the formation of the DL in haloes with mass close to the most efficient in producing DL, their contribution becomes almost negligible in very small haloes, but still important in Milky-Way like haloes, between 24% and 46%. It must be noted, however, that the different distributions are a natural consequence of the fact that larger haloes can host larger satellites, which are also those more subject to stripping. Once the x-axis in Figure 5 is scaled to the mass of the host (bottom panel), the main differences in the distributions disappear (see the discussion on single case studies in the next section) and the peak lies between $0.0030.001<\log(M_{gal}/M_{halo})<$ 0.003.
@@ 4 Discussion and Conclusion
The first theoretical prediction of the DL fraction in such small scales is dated to almost two decades ago, when Purcell et al. (2007) showed that their analytic model predicts an increasing fraction of DL for haloes in the range $1510.5<\log M_{halo}<$ 15, negligible below $\log M_{halo}\sim 12$, but around 30% on cluster scales. Their work gave the first hint for a significant different amount of DL in haloes much smaller than those usually probed in observations ($\log M_{halo}>12.5$). However, in the past few years there have been the first measurements of the DL in the Milky-Way (Deason et al. 2019) and M31 (Harmsen et al. 2017), and given the improvements in the spatial and mass resolutions of hydro-simulations, also the first theoretical predictions (e.g., Proctor et al. 2023; Ahvazi et al. 2023). Our model has never been tested on such small scales, and this Letter brings the first results on the formation of the DL in haloes below the typical halo masses usually probed.
The first key-point of this work is that the fraction of DL in Milky-Way like haloes is in good agreement with the observations of the stellar halo of the Milky-Way by Deason et al. (2019), and that of M31 by Harmsen et al. (2017) (Fig.1). Not only, our model predicts fractions that are in good agreement with the observations by Ragusa et al. (2023) and Kluge et al. (2021) in that halo mass range. Overall, the model agrees very well with the predictions of the Eagle project (Proctor et al. 2023) up to $\log M_{halo}\sim 12.1$, while those of TNG50 (Ahvazi et al. 2023) are biased-low with respect to ours, and predicts DL fractions that decrease quite importantly towards low mass haloes. The disagreement between Eagle and TNG50 simulations, in our opinion, has to be taken with caution. In fact, the two simulations have different mass resolutions and more importantly, the authors adopt different definitions for the DL. On one hand, Proctor et al. (2023) use Gaussian Mixture Models to separate bulge, disk and DL. On the other hand, Ahvazi et al. (2023) define DL the amount of stars that are bound only to the potential well of the main halo in the range $0.15R_{200}<r<R_{200}$.
@@
Most of the DL coming from stellar stripping, which represent almost the total, is stripped by intermediate/massive galaxies even in some haloes of the sample probed in this work. However, this contribution becomes less and less important when the halo mass decreases, and at Milky-Way/M31 like halo scales it accounts for around 30-35% (between 24% and 46%), to eventually drop to 16% at lower masses. The low-mass scales are characterized by haloes that produce their DL by stripping dwarf galaxies, with episodic events of stripping of more massive satellites. This has an important effect on the production of DL in such small haloes, because it will eventually become inefficient and the fractional budget of the DL is expected to drop to zero in halo masses lower than those probed here, as also predicted by the analytic model of Purcell et al. (2007).
@@
To summarize, the formation of the DL in haloes with mass lower than $\log M_{halo}=13$, which includes also Milky-Way like haloes, is driven by the halo concentration. More concentrated and relaxed haloes host a higher fraction of DL that forms from stellar stripping of satellite galaxies, with more massive satellites contributing mostly in the high mass end and dwarfs in the low mass end. However, on these scales most of the haloes are actually still on the way of their assembly, and so is the DL they host. It is reasonable to think that a thorough analysis on scales lower than those probed here would reveal a drop in the DL fraction, simply because the mechanism responsible for the DL formation is inhibited by the lack of sufficiently massive satellite galaxies.
@@ Acknowledgements
E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (2020R1A2C3003769). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (2022R1A6A1A03053472). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT. As the head of the group, S.K.Y. acted as a corresponding author.

### 29 Feb 2024 v2

# Diffuse Light in Milky Way-like Haloes

@@ Abstract
We investigate the diffuse light (DL) content of dark matter haloes in the mass range $11.5\leq\log M_{halo}\leq 13$, a range that includes also the dark matter halo of the Milky-Way, taking advantage of a state-of-the-art semi-analytic model run on the merger trees extracted from a set of high-resolution cosmological simulations. The fraction of DL in such relatively small haloes is found to progressively decrease from the high to the low mass end, in good agreement with analytic (Purcell et al. 2007) and numerical results from simulations (Proctor et al. 2023; Ahvazi et al. 2023), in good agreement also with the fraction of the DL observed in the Milky Way (Deason et al. 2019) and M31 (Harmsen et al. 2017). Haloes with different masses have a different efficiency in producing DL: $\log M_{halo}\simeq 13$ is found to be the characteristic halo mass where the production of DL is the most efficient, while the overall efficiency decreases at both larger (Contini et al. 2024) and smaller scales (this work). The DL content in this range of halo mass is the result of stellar stripping due to tidal interaction between satellites and its host (95%) and mergers between satellites and the central galaxy (5%), with pre-processed material, sub-channel of mergers and stripping and so already included in the 100%, that contributes no more than 8% on average. The halo concentration is the main driver of the DL formation: more concentrated haloes have higher DL fractions that come from stripping of more massive satellites in the high halo mass end, while dwarfs contribute mostly in the low halo mass end.
@@ 3 Results
Most of the DL forms through stellar stripping, and we already know (Contini et al. 2014) that intermediate/massive galaxies are responsible for the bulk of it. The last question we want to answer, in order to complete the analysis is: are intermediate/massive galaxies still the main contributors? The answer to this question is in Figure 5, where we plot the percentage of DL that comes from stellar stripping of galaxies having a given stellar mass (top panel), for haloes in different mass ranges as indicated in the legend of the Figure. We can clearly see that the peak of the distributions progressively decreases, from $\log M_{*}\sim 10.3$ in the highest halo mass range, to $\log M_{*}\sim 9.0$ in the lowest. The contribution of galaxies in the stellar mass range $11.09.5<\log M_{*}<$ 11.0 increases from 16% to 62% in the same ranges, implying that while intermediate/massive galaxies still contribute remarkably in the formation of the DL in haloes with mass close to the most efficient in producing DL, their contribution becomes almost negligible in very small haloes, but still important in Milky Way-like haloes, between 24% and 46%. It must be noted, however, that the different distributions are a natural consequence of the fact that larger haloes can host larger satellites, which are also those more subject to stripping. Once the x-axis in Figure 5 is scaled to the mass of the host (bottom panel), the main differences in the distributions disappear (see the discussion on single case studies in the next section) and the peak lies between $0.0030.001<\log(M_{gal}/M_{halo})<$ 0.003.
@@ 4 Discussion and Conclusion
The first theoretical prediction of the DL fraction in such small scales is dated to almost two decades ago, when Purcell et al. (2007) showed that their analytic model predicts an increasing fraction of DL for haloes in the range $1510.5<\log M_{halo}<$ 15, negligible below $\log M_{halo}\sim 12$, but around 30% on cluster scales. Their work gave the first hint for a significant different amount of DL in haloes much smaller than those usually probed in observations ($\log M_{halo}>12.5$). However, in the past few years there have been the first measurements of the DL in the Milky Way (Deason et al. 2019) and M31 (Harmsen et al. 2017), and given the improvements in the spatial and mass resolutions of hydro-simulations, also the first theoretical predictions (e.g., Proctor et al. 2023; Ahvazi et al. 2023). Our model has never been tested on such small scales, and this Letter brings the first results on the formation of the DL in haloes below the typical halo masses usually probed.
The first key-point of this work is that the fraction of DL in Milky Way-like haloes is in good agreement with the observations of the stellar halo of the Milky Way by Deason et al. (2019), and that of M31 by Harmsen et al. (2017) (Fig.1). Not only, our model predicts fractions that are in good agreement with the observations by Ragusa et al. (2023) and Kluge et al. (2021) in that halo mass range. Overall, the model agrees very well with the predictions of the Eagle project (Proctor et al. 2023) up to $\log M_{halo}\sim 12.1$, while those of TNG50 (Ahvazi et al. 2023) are biased-low with respect to ours, and predicts DL fractions that decrease quite importantly towards low mass haloes. The disagreement between Eagle and TNG50 simulations, in our opinion, has to be taken with caution. In fact, the two simulations have different mass resolutions and more importantly, the authors adopt different definitions for the DL. On one hand, Proctor et al. (2023) use Gaussian Mixture Models to separate bulge, disk and DL. On the other hand, Ahvazi et al. (2023) define DL the amount of stars that are bound only to the potential well of the main halo in the range $0.15R_{200}<r<R_{200}$.
@@
Most of the DL coming from stellar stripping, which represent almost the total, is stripped by intermediate/massive galaxies even in some haloes of the sample probed in this work. However, this contribution becomes less and less important when the halo mass decreases, and at Milky Way/M31 like halo scales it accounts for around 30-35% (between 24% and 46%), to eventually drop to 16% at lower masses. The low-mass scales are characterized by haloes that produce their DL by stripping dwarf galaxies, with episodic events of stripping of more massive satellites. This has an important effect on the production of DL in such small haloes, because it will eventually become inefficient and the fractional budget of the DL is expected to drop to zero in halo masses lower than those probed here, as also predicted by the analytic model of Purcell et al. (2007).
@@
To summarize, the formation of the DL in haloes with mass lower than $\log M_{halo}=13$, which includes also Milky Way-like haloes, is driven by the halo concentration. More concentrated and relaxed haloes host a higher fraction of DL that forms from stellar stripping of satellite galaxies, with more massive satellites contributing mostly in the high mass end and dwarfs in the low mass end. However, on these scales most of the haloes are actually still on the way of their assembly, and so is the DL they host. It is reasonable to think that a thorough analysis on scales lower than those probed here would reveal a drop in the DL fraction, simply because the mechanism responsible for the DL formation is inhibited by the lack of sufficiently massive satellite galaxies.
@@ Acknowledgements
E.C. and S.K.Y. acknowledge support from the Korean National Research Foundation (2020R1A2C3003769). E.C. and S.J. acknowledge support from the Korean National Research Foundation (RS-2023-00241934). All the authors are supported by the Korean National Research Foundation (2022R1A6A1A03053472). J.R. was supported by the KASI-Yonsei Postdoctoral Fellowship and by the Korea Astronomy and Space Science Institute under the R&D program (Project No. 2023-1-830-00), supervised by the Ministry of Science and ICT.